# High Dimensional Distributed Gradient Descent with Arbitrary Number of Byzantine Attackers

### v2

@@ A.1 Proof of Lemma 3
Denote $\lambda_{k+1}(S)$ as the $(k+1)$-th largest eigenvalue of $V(S)$. Recall step 1 and 1 in Algorithm 1, $V(S)=\sum_{i=1}^{d}\lambda_{i}(S)\mathbf{u}_{i}\mathbf{u}_{i}^{T}$, in which $\lambda_{i}(S)$ is the $i$-th largest eigenvalue of $V(S)$. Therefore
$(\mathbf{I}-\mathbf{P})V(S)(\mathbf{I}-\mathbf{P})=\sum_{i=k+1}^{d}\lambda_{i}(S)\mathbf{u}_{i}\mathbf{u}_{i}^{T},$ (57)
hence for any $\mathbf{u}$ with $\|\mathbf{u}\|_{2}=1$, $\mathbf{u}^{T}(\mathbf{I}-\mathbf{P})V(S)(\mathbf{I}-\mathbf{P})\mathbf{u}\leq\lambda_{k+1}(S)$. Let
@@
It remains to bound $|S|/|G|$. From (10) and (11), From now on, suppose the following condition hold: We will show that (47) and (48) hold with high probability later. Now suppose that these two conditions hold for all iterations, then the following lemma holds:
@@
The proof is shown in section A.2. ∎
@@ A.2 Proof of Lemma 6
The proof is similar to the proof of Lemma 3 in section A.1. Denote $Q_{t}=S_{t}\setminus G_{t}$. For any $\mathbf{u}$ with $\|\mathbf{u}\|_{2}=1$,
@@
$\frac{1}{|G_{t}|}\sum_{i\in G}\tau_{i}$$\leq$ $8\sigma^{2}\left(1+\frac{2d}{\alpha N}\right)\frac{p}{\lambda_{p}(S_{t})}+\frac{|S_{t}|}{|G_{t}|},$ (69)
@@
$\overset{(a)}{\leq}$ $8\sigma^{2}\left(1+\frac{2d}{\alpha N}\right)\operatorname{tr}(\mathbf{P}V^{-1}(S_{t})\mathbf{P})+\|PV^{-\frac{1}{2}}(S_{t})(\mu(G_{t})-\mu(S_{t}))\|_{2}^{2}$
$\overset{(b)}{\leq}$ $8\sigma^{2}\left(1+\frac{2d}{\alpha N}\right)\operatorname{tr}(\mathbf{P}V^{-1}(S_{t})\mathbf{P})+\frac{|S_{t}|}{|G_{t}|}$
$\overset{(c)}{\leq}$ $8\sigma^{2}\left(1+\frac{2d}{\alpha N}\right)\sum_{i=1}^{m}\frac{1}{\lambda_{i}(S_{t})}+\frac{|S_{t}|}{|G_{t}|}$
$\leq$ $8\sigma^{2}\left(1+\frac{2d}{\alpha N}\right)\frac{p}{\lambda_{p}(S_{t})}+\frac{|S_{t}|}{|G_{t}|}.$
@@
$\frac{1}{|G_{t}|}\sum_{i\in G}(\mathbf{Y}_{i}-\mu(G_{t}))(\mathbf{Y}_{i}-\mu(G_{t}))^{T}\preceq\frac{1}{|G_{t}|}\sum_{i\in G}(\mathbf{Y}_{i}-\mu^{*})(\mathbf{Y}_{i}-\mu^{*})^{T}\leq 8\sigma^{2}\left(1+\frac{2d}{\alpha N}\right).$ (72)
@@
$V(S_{t})=\sum_{i=1}^{d}\lambda_{i}(S_{t})\mathbf{u}_{i}\mathbf{u}_{i}^{T},$ (73)
and $\mathbf{P}\mathbf{u}_{i}=\mathbf{u}_{i}$ for $i\leq m$, $\mathbf{P}\mathbf{u}_{i}=\mathbf{0}$ for $i>m$, then
$\mathbf{P}V^{-1}(S_{t})\mathbf{P}=\sum_{i=1}^{d}\frac{1}{\lambda_{i}(S_{t})}\mathbf{P}\mathbf{u}_{i}\mathbf{u}_{i}^{T}\mathbf{P}=\sum_{i=1}^{m}\frac{1}{\lambda_{i}(S_{t})}\mathbf{u}_{i}\mathbf{u}_{i}^{T},$ (74)
thus $\operatorname{tr}(\mathbf{P}V^{-1}(S_{t})\mathbf{P})=\sum_{i=1}^{m}1/\lambda_{i}(S_{t})$.
@@
$\lambda_{p}(S_{t})\geq\lambda_{c}\geq 32\sigma^{2}\left(1+\frac{2d}{\alpha N}\right)\frac{1+\epsilon}{1-\epsilon},$ (76)
@@
$\frac{1}{|G_{t}|}\sum_{i\in G}\tau_{i}$$\leq$ $8\sigma^{2}\left(1+\frac{2d}{\alpha N}\right)\frac{p}{\lambda_{p}(S_{t})}+\frac{|S_{t}|}{|G_{t}|}$
@@
A.3 Proof of Lemma 4
Suppose there are $n$ independent random variables $Z_{1},\ldots,Z_{n}$. $\text{P}(Z_{i}=1)=p_{i}$, and $\text{P}(Z_{i}=0)=1-p_{i}$. Let $\bar{p}=(1/n)\sum_{i=1}^{n}p_{i}$, then we show a concentration inequality of the mean of $Z_{i}$.
Let $Z=\sum_{i=1}^{n}Z_{i}$. Then
$\mathbb{E}[e^{\lambda Z}]=\mathbb{E}\left[\exp\left(\lambda\sum_{i=1}^{n}Z_{i}\right)\right]=\Pi_{i=1}^{n}\mathbb{E}[e^{\lambda Z_{i}}]=\Pi_{i=1}^{n}(1-p_{i}+p_{i}e^{\lambda})$ (82)
$\leq$ $\exp\left(\sum_{i=1}^{n}\ln(1-p_{i}+p_{i}e^{\lambda})\right)\leq\exp\left(\sum_{i=1}^{n}p_{i}(e^{\lambda}-1)\right)=\exp\left(n\bar{p}(e^{\lambda}-1)\right).$
@@ A.3 Proof of Lemma 4
$\text{P}(Z>n\bar{p}(1+\epsilon))$$\leq$ $\underset{\lambda\geq 0}{\inf}\exp\left[-n\bar{p}(1+\epsilon)\lambda+n\bar{p}(e^{\lambda}-1)\right]$ (83)
$=$ $\exp\left[-n\bar{p}((1+\epsilon)\ln(1+\epsilon)-\epsilon)\right].$
@@
$\text{P}(Z<n\bar{p}(1-\epsilon))$$\leq$ $\underset{\lambda\geq 0}{\inf}\exp\left[n\bar{p}(1+\epsilon)\lambda+n\bar{p}(e^{-\lambda}-1)\right]$ (84)
$=$ $\exp\left[-n\bar{p}((1-\epsilon)\ln(1-\epsilon)-\epsilon)\right].$
@@
$\text{P}\left(\left|\frac{1}{n\bar{p}}\sum_{i=1}^{n}Z_{i}-1\right|>\epsilon\right)\leq 2e^{-\frac{1}{4}n\bar{p}\epsilon^{2}}.$ (85)
@@
$\bar{p}\overset{(a)}{\geq}\frac{1}{|S_{t}|}\sum_{i\in S}\frac{\tau_{i}}{\tau_{\max}}\overset{(b)}{=}\frac{p}{\tau_{\max}}\overset{(c)}{\geq}\frac{\lambda_{c}m}{2N^{\frac{2}{3}}},$ (86)
in which (a) comes from step 1 in Algorithm 1. (b) comes from (70). (c) holds because the prefilter step (step 1) in Algorithm 1, as well as (6). Hence, by taking union bound,
$\text{P}\left(\exists t,\frac{|L_{Gt}|}{|G_{t}|}>\frac{1+\epsilon}{|G_{t}|}\sum_{i\in G}\frac{\tau_{i}}{\tau_{\max}}\text{ or }\frac{|L_{t}|}{|S_{t}|}<\frac{1-\epsilon}{|S_{t}|}\sum_{i\in S}\frac{\tau_{i}}{\tau_{\max}}\right)\leq 4N\exp\left[-\frac{1}{16}\lambda_{c}m\alpha^{2}N^{\frac{1}{3}}\epsilon^{2}\right].$ (87)
A.4 Proof of Lemma 5
@@
=$\overset{(a)}{=}$ $\underset{\|\mathbf{u}\|_{2}=1}{\max}\mathbf{u}^{T}\frac{1}{|G|}\sum_{i\in G}(\mathbf{X}_{i}-\mu^{*})(\mathbf{X}_{i}-\mu^{*})^{T}\mathbf{u}$
@@
$\overset{(b)}{\leq}$ $8\sigma^{2}\left(1+\frac{2d}{\alpha N}\right)\frac{|G_{0}|}{|G|}$
<$\overset{(c)}{<}$ $\frac{\lambda_{c}}{2\alpha},$
in which (a) comes from the definition that $G_{0}\subset S_{0}^{*}$. Therefore, $\mathbf{Y}_{i}$ is not modified from the clean sample $\mathbf{X}_{i}$. (b) use Lemma 1. (c) uses $\lambda_{c}>32\sigma^{2}(1+2d/(\alpha N))$, $|G_{0}|=\alpha N/2$, and from Corollary 1, $|G|\geq\alpha^{2}N/4$.
@@ Appendix B Proof of Theorem 2
Our proof begins with the matrix Bernstein inequality [46].
@@
(From [46], Theorem 1.6.2) Let $\mathbf{W}_{1},\ldots,\mathbf{W}_{N}$ be i.i.d random matrices, with $\mathbb{E}[\mathbf{W}_{i}]=\mathbf{0}$, and $\|\mathbf{W}_{i}\|\leq L$ almost surely. Define $\mathbf{Z}=\sum_{i=1}^{N}\mathbf{W}_{i}$. Then for all $t\geq 0$,
@@
With probability at least $1-\exp[-((1/2)\ln 2-1/4)N\alpha]-2d\exp(-(\ln^{2}N)/2)$, there exists a set $G_{0}\subset S_{0}^{*}$ with $|G_{0}|\geq\alpha N/2$, and
$\left\|\frac{1}{|G_{0}|}\sum_{i\in G_{0}}(\mathbf{Y}_{i}-\mu^{*})(\mathbf{Y}_{i}-\mu^{*})^{T}\right\|_{op}\leq\frac{2\sigma^{2}}{\alpha}\left(1+\sqrt{\frac{16d\ln^{2}N}{3\alpha N}}\right)^{2}.$ (91)
@@
Recall that $\mathbf{X}_{i}$ for all $i\in S_{0}$ are samples before contamination, while $\mathbf{Y}_{i}$ are observed values. $\mathbf{Y}_{i}=\mathbf{X}_{i}$ for at least $\alpha N$ samples.
@@
$\mathbb{E}[\mathbf{W}\mathbf{W}^{T}]$ $=$ $\frac{1}{N^{2}}\mathbb{E}\left[(\mathbf{X}-\mu^{*})(\mathbf{X}-\mu^{*})^{T}(\mathbf{X}-\mu^{*})(\mathbf{X}-\mu^{*})^{T}\mathbf{1}(\|\mathbf{X}-\mu^{*}\|\leq M)\right]$ (94)
@@
$\preceq$ $\frac{M^{2}}{N^{2}}\mathbb{E}\left[(\mathbf{X}-\mu^{*})(\mathbf{X}-\mu^{*})^{T}\right]$
$=$ $\frac{M^{2}}{N^{2}}V^{*}\preceq\frac{M^{2}\sigma^{2}}{N^{2}}\mathbf{I}.$
@@
$\text{P}\left(n>\frac{1}{2}\alpha N\right)\leq\exp\left[-\left(\frac{1}{2}\ln 2-\frac{1}{4}\right)N\alpha\right].$ (101)
Now we assume that $\|\mathbf{Z}\|\leq t_{0}$ and $n\leq\alpha N/2$. The probability of violating these two conditions are bounded in (97) and (101), respectively. Then define
@@
$|G_{0}|\geq N-|\mathcal{B}|-n\geq N-(1-\alpha)N-\frac{1}{2}\alpha N=\frac{1}{2}\alpha N.$ (103)
@@
=$\overset{(a)}{=}$ $\left\|\frac{1}{|G_{0}|}\sum_{i\in G_{0}}(\mathbf{X}_{i}-\mu^{*})(\mathbf{X}_{i}-\mu^{*})^{T}\right\|$
@@
=$\overset{(c)}{=}$ $\frac{2}{\alpha}\sigma^{2}\left(1+\frac{M}{\sigma}\sqrt{\frac{2\ln^{2}N}{N}}+\frac{4M^{2}}{3N\sigma^{2}}\ln^{2}N\right)$
=$\overset{(d)}{=}$ $\frac{2}{\alpha}\sigma^{2}\left(1+\sqrt{\frac{8d\ln^{2}N}{\alpha N}}+\frac{16d}{3\alpha}\frac{\ln^{2}N}{N}\right)$
$<$ $\frac{2\sigma^{2}}{\alpha}\left(1+\sqrt{\frac{16d\ln^{2}N}{3\alpha N}}\right)^{2}.$
For (a), recall that for all samples that are not attacked, $\mathbf{Y}_{i}=\mathbf{X}_{i}$. (b) comes from (98). (c) comes from (96). (d) comes from (99). Recall that (104) relies on two conditions, $\|\mathbf{Z}\|\leq t_{0}$ and $n\leq\alpha N/2$. Therefore, the probability that (104) does not hold is no more than the sum of (97) and (101). The proof is complete. ∎
@@
$\lambda_{c}\geq\frac{8\sigma^{2}}{\alpha}\left(1+\sqrt{\frac{16d\ln^{2}N}{3\alpha N}}\right)^{2},$ (105)
@@
Appendix C Proof of Theorem 3
@@
With probability at least $1-\exp[-(\ln 2-1/2)N\alpha]$,
@@
It remains to get a lower bound of $R_{A}^{\prime}(\beta)$. Let $Z$ be a random variable taking values in $\{1,\ldots,d\}$ uniformly, with each value of $Z$ corresponds to a distribution of $\mathbf{X}$. Denote $p_{i}(\cdot)=\mathbf{P}(\cdot|Z=i)$. Under $Z=i$.
@@
Construct $d$ hypotheses, and denote $p_{i}$ as the distribution of $\mathbf{X}$ under $i$-th hypothesis. If $Z=i$, then $\mathbf{X}$ follows $p_{i}$. $p_{i}$ is defined as follows. Given $Z=i$, $X_{i}=\sigma/\sqrt{2\beta}$. For all $j\neq i$, $p_{i}(X_{j}=0)=1-2\beta$, $p_{i}(X_{j}=\sigma/\sqrt{2\beta})=\beta$, and $p_{i}(X_{j}=-\sigma/\sqrt{2\beta})=\beta$. Conditional on $Z=i$, $\mathbf{X}_{j}$ are independent for $j\in[d]\setminus\{i\}$.
The above construction uses the idea from [37], Proposition 5.4.
The mean of $p_{i}$ is
$\mu_{i}=\frac{\sigma}{\sqrt{2\beta}}\mathbf{e}_{i},$ (111)
@@
Construct $p_{Y}$, the distribution of $Y$ as following: $p_{Y}(Y_{j}=0)=1-2\beta$, $p_{Y}(Y_{j}=\sigma/\sqrt{2\beta})=p_{Y}(Y_{j}=-\sigma/\sqrt{2\beta})=\beta$. Then it is straightforward to show that for each $i$, there exists a $q_{i}$ such that $p_{Y}=\beta p_{i}+(1-\beta)q_{i}$. Therefore, we design the contamination strategy $\pi_{A}^{\prime}(\beta)$ as follows: with probability $\beta$, $\mathbf{Y}=\mathbf{X}$; with probability $1-\beta$, the sample is corrupted, and then $\mathbf{Y}$ follows a conditional distribution $q_{i}$. With this operation, for all values of $Z$, the distribution of $\mathbf{Y}$ is always $p_{Y}$, which does not depend on $Z$.

### v3

@@ Appendix A Appendix
Recall Algorithm 2, $V(S)=\sum_{j=1}^{d}\lambda_{j}(S)\mathbf{u}_{j}\mathbf{u}_{j}^{T}$, in which $\lambda_{j}(S)$ is the $j$-th largest eigenvalue of $V(S)$. Therefore
$(\mathbf{I}-\mathbf{P})V(S)(\mathbf{I}-\mathbf{P})=\sum_{j=p+1}^{d}\lambda_{j}(S)\mathbf{u}_{j}\mathbf{u}_{j}^{T},$ (55)
hence for any $\mathbf{u}$ with $\|\mathbf{u}\|_{2}=1$, $\mathbf{u}^{T}(\mathbf{I}-\mathbf{P})V(S)(\mathbf{I}-\mathbf{P})\mathbf{u}\leq\lambda_{p+1}(S)$. Let
@@
It remains to bound $|S|/|G|$.
@@
The proof is shown in section A. ∎
@@
The proof is similar to the proof of Lemma 3 in section A. Denote $Q_{t}=S_{t}\setminus G_{t}$. For any $\mathbf{u}$ with $\|\mathbf{u}\|_{2}=1$,
@@
$\frac{1}{|G_{t}|}\sum_{i\in G}\tau_{i}$ $\leq$ $8\sigma^{2}\left(1+\frac{2d}{\alpha m}\right)\frac{p}{\lambda_{p}(S_{t})}+\frac{|S_{t}|}{|G_{t}|},$ (67)
@@
$\overset{(a)}{\leq}$ $8\sigma^{2}\left(1+\frac{2d}{\alpha m}\right)\operatorname{tr}(\mathbf{P}V^{-1}(S_{t})\mathbf{P})+\|PV^{-\frac{1}{2}}(S_{t})(\mu(G_{t})-\mu(S_{t}))\|_{2}^{2}$
$\overset{(b)}{\leq}$ $8\sigma^{2}\left(1+\frac{2d}{\alpha m}\right)\operatorname{tr}(\mathbf{P}V^{-1}(S_{t})\mathbf{P})+\frac{|S_{t}|}{|G_{t}|}$
$\overset{(c)}{\leq}$ $8\sigma^{2}\left(1+\frac{2d}{\alpha m}\right)\sum_{j=1}^{m}\frac{1}{\lambda_{j}(S_{t})}+\frac{|S_{t}|}{|G_{t}|}$
$\leq$ $8\sigma^{2}\left(1+\frac{2d}{\alpha m}\right)\frac{p}{\lambda_{p}(S_{t})}+\frac{|S_{t}|}{|G_{t}|}.$
@@
$\frac{1}{|G_{t}|}\sum_{i\in G}(\mathbf{Y}_{i}-\mu(G_{t}))(\mathbf{Y}_{i}-\mu(G_{t}))^{T}\preceq\frac{1}{|G_{t}|}\sum_{i\in G}(\mathbf{Y}_{i}-\mu^{*})(\mathbf{Y}_{i}-\mu^{*})^{T}\leq 8\sigma^{2}\left(1+\frac{2d}{\alpha m}\right).$ (70)
@@
$V(S_{t})=\sum_{j=1}^{d}\lambda_{j}(S_{t})\mathbf{u}_{j}\mathbf{u}_{j}^{T},$ (71)
and $\mathbf{P}\mathbf{u}_{j}=\mathbf{u}_{j}$ for $i\leq m$, $\mathbf{P}\mathbf{u}_{j}=\mathbf{0}$ for $i>m$, then
$\mathbf{P}V^{-1}(S_{t})\mathbf{P}=\sum_{j=1}^{d}\frac{1}{\lambda_{j}(S_{t})}\mathbf{P}\mathbf{u}_{j}\mathbf{u}_{j}^{T}\mathbf{P}=\sum_{i=1}^{m}\frac{1}{\lambda_{j}(S_{t})}\mathbf{u}_{j}\mathbf{u}_{j}^{T},$ (72)
thus $\operatorname{tr}(\mathbf{P}V^{-1}(S_{t})\mathbf{P})=\sum_{i=1}^{m}1/\lambda_{j}(S_{t})$.
@@
$\lambda_{p}(S_{t})\geq\lambda_{c}\geq 32\sigma^{2}\left(1+\frac{2d}{\alpha m}\right)\frac{1+\epsilon}{1-\epsilon},$ (74)
@@
$\frac{1}{|G_{t}|}\sum_{i\in G}\tau_{i}$ $\leq$ $8\sigma^{2}\left(1+\frac{2d}{\alpha m}\right)\frac{p}{\lambda_{p}(S_{t})}+\frac{|S_{t}|}{|G_{t}|}$
@@
Proof of Lemma 4
Suppose there are $k$ independent random variables $Z_{1},\ldots,Z_{k}$. $\text{P}(Z_{i}=1)=p_{i}$, and $\text{P}(Z_{i}=0)=1-p_{i}$. Let $\bar{p}=(1/k)\sum_{i=1}^{n}p_{i}$, then we show a concentration inequality of the mean of $Z_{i}$.
Let $Z=\sum_{i=1}^{k}Z_{i}$. Then
$\mathbb{E}[e^{\lambda Z}]=\mathbb{E}\left[\exp\left(\lambda\sum_{i=1}^{k}Z_{i}\right)\right]=\Pi_{i=1}^{k}\mathbb{E}[e^{\lambda Z_{i}}]=\Pi_{i=1}^{k}(1-p_{i}+p_{i}e^{\lambda})$ (80)
$\leq$ $\exp\left(\sum_{i=1}^{n}\ln(1-p_{i}+p_{i}e^{\lambda})\right)\leq\exp\left(\sum_{i=1}^{k}p_{i}(e^{\lambda}-1)\right)=\exp\left(k\bar{p}(e^{\lambda}-1)\right).$
@@
$\text{P}(Z>k\bar{p}(1+\epsilon))$ $\leq$ $\underset{\lambda\geq 0}{\inf}\exp\left[-k\bar{p}(1+\epsilon)\lambda+k\bar{p}(e^{\lambda}-1)\right]$ (81)
$=$ $\exp\left[-k\bar{p}((1+\epsilon)\ln(1+\epsilon)-\epsilon)\right].$
@@
$\text{P}(Z<k\bar{p}(1-\epsilon))$ $\leq$ $\underset{\lambda\geq 0}{\inf}\exp\left[k\bar{p}(1+\epsilon)\lambda+k\bar{p}(e^{-\lambda}-1)\right]$ (82)
$=$ $\exp\left[-k\bar{p}((1-\epsilon)\ln(1-\epsilon)-\epsilon)\right].$
@@
$\text{P}\left(\left|\frac{1}{k\bar{p}}\sum_{i=1}^{n}Z_{i}-1\right|>\epsilon\right)\leq 2e^{-\frac{1}{4}k\bar{p}\epsilon^{2}}.$ (83)
@@
$\bar{p}\overset{(a)}{\geq}\frac{1}{|S_{t}|}\sum_{i\in S}\frac{\tau_{i}}{\tau_{\max}}\overset{(b)}{=}\frac{p}{\tau_{\max}}\overset{(c)}{\geq}\frac{\lambda_{c}}{m^{\frac{2}{3}}},$ (84)
in which (a) comes from Algorithm 2. (b) comes from (68). For (c),
$\tau_{i}$ $=$ $\|\mathbf{P}V^{-\frac{1}{2}}(s)(\mathbf{Y}_{i}-\mu(S))\|_{2}^{2}$ (85)
$=$ $\|\sum_{j=1}^{p}\lambda_{j}^{-\frac{1}{2}}\mathbf{u}_{j}\mathbf{u}_{j}^{T}(\mathbf{Y}_{j}-\mu(S))\|_{2}^{2}$
$\leq$ $p\lambda_{c}^{-1}\|\mathbf{Y}_{j}-\mu(S)\|_{2}^{2}$
$\leq$ $p\lambda_{c}^{-1}m^{\frac{2}{3}}.$
Hence, by taking union bound, and use $k\geq\alpha m/2$, we have
$\text{P}\left(\exists t,\frac{|L_{Gt}|}{|G_{t}|}>\frac{1+\epsilon}{|G_{t}|}\sum_{i\in G}\frac{\tau_{i}}{\tau_{\max}}\text{ or }\frac{|L_{t}|}{|S_{t}|}<\frac{1-\epsilon}{|S_{t}|}\sum_{i\in S}\frac{\tau_{i}}{\tau_{\max}}\right)\leq 4m\exp\left[-\frac{1}{8}\alpha\lambda_{c}m^{\frac{1}{3}}\epsilon^{2}\right].$ (86)
Proof of Lemma 5
@@
$\overset{(a)}{=}$ $\underset{\|\mathbf{u}\|_{2}=1}{\max}\mathbf{u}^{T}\frac{1}{|G|}\sum_{i\in G}(\mathbf{X}_{i}-\mu^{*})(\mathbf{X}_{i}-\mu^{*})^{T}\mathbf{u}$
@@
$\overset{(b)}{\leq}$ $8\sigma^{2}\left(1+\frac{2d}{\alpha m}\right)\frac{|G_{0}|}{|G|}$
$\overset{(c)}{<}$ $\frac{\lambda_{c}}{2\alpha},$
in which (a) comes from the definition that $G_{0}\subset S_{0}^{*}$. Therefore, $\mathbf{Y}_{i}$ is not modified from the clean sample $\mathbf{X}_{i}$. (b) use Lemma 1. (c) uses $\lambda_{c}>32\sigma^{2}(1+2d/(\alpha m))$, $|G_{0}|=\alpha m/2$, and from Corollary 1, $|G|\geq\alpha^{2}N/4$.
@@
Our proof begins with the matrix Bernstein inequality (tropp2015introduction).
@@
(From (tropp2015introduction), Theorem 1.6.2) Let $\mathbf{W}_{1},\ldots,\mathbf{W}_{N}$ be i.i.d random matrices, with $\mathbb{E}[\mathbf{W}_{i}]=\mathbf{0}$, and $\|\mathbf{W}_{i}\|\leq L$ almost surely. Define $\mathbf{Z}=\sum_{i=1}^{N}\mathbf{W}_{i}$. Then for all $t\geq 0$,
@@
With probability at least $1-\exp[-((1/2)\ln 2-1/4)N\alpha]-2d\exp(-(\ln^{2}N)/2)$, there exists a set $G_{0}\subset S_{0}^{*}$ with $|G_{0}|\geq\alpha m/2$, and
$\left\|\frac{1}{|G_{0}|}\sum_{i\in G_{0}}(\mathbf{Y}_{i}-\mu^{*})(\mathbf{Y}_{i}-\mu^{*})^{T}\right\|_{op}\leq\frac{2\sigma^{2}}{\alpha}\left(1+\sqrt{\frac{16d\ln^{2}N}{3\alpha m}}\right)^{2}.$ (90)
@@
Recall that $\mathbf{X}_{i}$ for all $i\in S_{0}$ are samples before contamination, while $\mathbf{Y}_{i}$ are observed values. $\mathbf{Y}_{i}=\mathbf{X}_{i}$ for at least $\alpha m$ samples.
@@
$\mathbb{E}[\mathbf{W}\mathbf{W}^{T}]$ $=$ $\frac{1}{m^{2}}\mathbb{E}\left[(\mathbf{X}-\mu^{*})(\mathbf{X}-\mu^{*})^{T}(\mathbf{X}-\mu^{*})(\mathbf{X}-\mu^{*})^{T}\mathbf{1}(\|\mathbf{X}-\mu^{*}\|\leq M)\right]$ (93)
@@
$\preceq$ $\frac{M^{2}}{m^{2}}\mathbb{E}\left[(\mathbf{X}-\mu^{*})(\mathbf{X}-\mu^{*})^{T}\right]$
$=$ $\frac{M^{2}}{m^{2}}V^{*}\preceq\frac{M^{2}\sigma^{2}}{m^{2}}\mathbf{I}.$
@@
$\text{P}\left(n>\frac{1}{2}\alpha m\right)\leq\exp\left[-\left(\frac{1}{2}\ln 2-\frac{1}{4}\right)N\alpha\right].$ (100)
Now we assume that $\|\mathbf{Z}\|\leq t_{0}$ and $n\leq\alpha m/2$. The probability of violating these two conditions are bounded in (96) and (100), respectively. Then define
@@
$|G_{0}|\geq N-|\mathcal{B}|-n\geq N-(1-\alpha)N-\frac{1}{2}\alpha m=\frac{1}{2}\alpha m.$ (102)
@@
$\overset{(a)}{=}$ $\left\|\frac{1}{|G_{0}|}\sum_{i\in G_{0}}(\mathbf{X}_{i}-\mu^{*})(\mathbf{X}_{i}-\mu^{*})^{T}\right\|$
@@
$\overset{(c)}{=}$ $\frac{2}{\alpha}\sigma^{2}\left(1+\frac{M}{\sigma}\sqrt{\frac{2\ln^{2}N}{N}}+\frac{4M^{2}}{3N\sigma^{2}}\ln^{2}N\right)$
$\overset{(d)}{=}$ $\frac{2}{\alpha}\sigma^{2}\left(1+\sqrt{\frac{8d\ln^{2}N}{\alpha m}}+\frac{16d}{3\alpha}\frac{\ln^{2}N}{N}\right)$
$<$ $\frac{2\sigma^{2}}{\alpha}\left(1+\sqrt{\frac{16d\ln^{2}N}{3\alpha m}}\right)^{2}.$
For (a), recall that for all samples that are not attacked, $\mathbf{Y}_{i}=\mathbf{X}_{i}$. (b) comes from (97). (c) comes from (95). (d) comes from (98). Recall that (103) relies on two conditions, $\|\mathbf{Z}\|\leq t_{0}$ and $n\leq\alpha m/2$. Therefore, the probability that (103) does not hold is no more than the sum of (96) and (100). The proof is complete. ∎
@@
$\lambda_{c}\geq\frac{8\sigma^{2}}{\alpha}\left(1+\sqrt{\frac{16d\ln^{2}N}{3\alpha m}}\right)^{2},$ (104)
@@
Proof of Theorem 3
Recall that $R_{A}$ and $R_{S}$ are defined as
and
in which $\mathcal{F}$ is the set of all distributions satisfying
$\operatorname{Cov}[\mathbf{X}_{i}]\preceq\frac{\sigma^{2}}{n}\mathbf{I},$ (108)
and
$\operatorname{Cov}[\mathbf{X}_{0}]\preceq\frac{\sigma^{2}}{N_{A}}\mathbf{I}.$ (110)
Under the condition
$\left\lceil\frac{N_{A}}{n}\right\rceil\leq\frac{\ln\frac{d}{4}}{4\beta\left(\ln\frac{1}{\beta}+1\right)},$ (111)
we need to prove
$R_{A}(\alpha)$ $\geq$ $\frac{\sigma}{2\sqrt{2\alpha}}\left(\frac{1}{\sqrt{n}}+\frac{1}{\sqrt{N_{A}}}\right),$ (112)
$R_{S}(\alpha)$ $\geq$ $\frac{\sigma}{2\sqrt{2}\alpha}\left(\frac{1}{\sqrt{n}}+\frac{1}{\sqrt{N_{A}}}\right).$ (113)
@@
With probability at least $1-\exp[-(\ln 2-1/2)m\alpha]$,
@@
It remains to get a lower bound of $R_{A}^{\prime}(\beta)$.
We analyze the case with $N_{A}\geq n$ first. Let
$k=\left\lceil\frac{N_{A}}{n}\right\rceil.$ (118)
Let $Z$ be a random variable taking values in $\{1,\ldots,d\}$ uniformly, with each value of $Z$ corresponds to a hypothesis. Denote $p_{j}(\cdot)$ as the distribution of the underlying truth of corrupted samples $\mathbf{X}_{1},\ldots,\mathbf{X}_{m}$ and the auxiliary clean sample $\mathbf{X}_{0}$ as following.
@@
Construct $d$ hypotheses. For $j$-th hypothesis, define a distribution $p_{j}$ as following: For $i=1,\ldots,m$, let $X_{ij}=\sigma/\sqrt{2n\beta}$, in which $X_{ij}$ is the $j$-th component of $\mathbf{X}_{i}$. For all $l\neq j$, $p_{i}(X_{il}=0)=1-2\beta$, $p_{i}(X_{il}=\sigma/\sqrt{2n\beta})=\beta$, and $p_{i}(X_{il}=-\sigma/\sqrt{2n\beta})=\beta$. Moreover, construct $k$ vectors $\mathbf{U}_{1},\ldots,\mathbf{U}_{k}$. Conditional on $Z=j$, $\mathbf{X}_{1},\ldots,\mathbf{X}_{m},\mathbf{U}_{1},\ldots,\mathbf{U}_{k}$ are i.i.d. Then let
$\mathbf{X}_{0}=\frac{1}{k}\sum_{i=1}^{k}\mathbf{U}_{i}.$ (119)
The above construction uses some idea from (diakonikolas2018list), Proposition 5.4.
The mean of $p_{j}$ is
$\mu_{j}=\frac{\sigma}{\sqrt{2n\beta}}\mathbf{e}_{j},$ (120)
@@
Construct $p_{Y}$, the distribution of $Y$ as following: $p_{Y}(Y_{j}=0)=1-2\beta$, $p_{Y}(Y_{j}=\sigma/\sqrt{2n\beta})=p_{Y}(Y_{j}=-\sigma/\sqrt{2n\beta})=\beta$. Then it is straightforward to show that for each $i$, there exists a $q_{i}$ such that $p_{Y}=\beta p_{i}+(1-\beta)q_{i}$. Therefore, we design the contamination strategy $\pi_{A}^{\prime}(\beta)$ as follows: for each $i\in[m]$, with probability $\beta$, $\mathbf{Y}_{i}=\mathbf{X}_{i}$; with probability $1-\beta$, the sample is corrupted, and then $\mathbf{Y}_{i}$ follows a conditional distribution $q_{i}$. With this operation, for all values of $Z$, the distribution of $\mathbf{Y}_{i}$ is always $p_{Y}$, which does not depend on $Z$.